# Online Cascade Learning for Efficient Inference over Streams

### v2

@@ 4 Experimental Setup
In our experiments, we construct a cascade system using three models: (i) a logistic regression model, (ii) a pretrained BERT-base model with 110M parameters (Kenton & Toutanova, 2019), and (iii) GPT-3.5 Turbo [^1] https://platform.openai.com/docs/models/gpt-3-5. This diverse set of models allows us to evaluate the effectiveness and robustness of our online cascade learning approach across various complexities and types of queries.
@@
4.1 Benchmarks
We select benchmarks that reflect the demands of real-world streaming applications. These tasks are integral to a wide range of commercial services, from customer feedback analysis to content moderation, thus offering a comprehensive testbed to evaluate our framework’s adaptability and efficiency in various practical scenarios.
IMDB
A binary sentiment classification benchmark with 50,000 movie reviews (Maas et al., 2011). The dataset has an even distribution of positive and negative samples. We utilize the official training split for our experiments, which contains 25,000 samples.
HateSpeech
A binary classification dataset consists of posts from an online forum, focusing on distinguishing between hate and noHate labels (de Gibert et al., 2018). After filtering, the dataset contains 10,703 samples with a pronounced class imbalance (1:7.95 ratio) between hatespeech and non-hatespeech examples. This imbalance mirrors a realistic challenge in streaming data environments, particularly in detecting harmful content. Our evaluation on this benchmark focuses on both accuracy and recall rate.
ISEAR
A multi-class emotion detection benchmark encompassing 7,666 samples across seven categories (Joy, Fear, Anger, Sadness, Disgust, Shame, Guilt) (Shao et al., 2015). Each category is well-represented, providing a balanced label distribution across the dataset.
FEVER
A fact-checking dataset with 6,512 claims manually verified against Wikipedia, labeled as Supported or Refuted (Thorne et al., 2018). It tests our framework’s ability to perform complex reasoning and information verification, a crucial aspect for real-time truth assessment in streaming data applications.
@@ 4.1 Benchmarks
4.2 Baselines
We compare our online cascade learning framework against several baselines to establish its effectiveness:
@@ 4.2 Baselines
  • Online Ensemble Learning: We employ all available models in an ensemble with predetermined probabilities. The smaller models are also continuously updated based on LLM annotations. This serves as an ablation of our online cascade learning by excluding the deferral policy learning component.
@@ 5.1 Overall Performance and Cost Trade-offs
The results on the HateSpeech dataset further reveal the strengths of our online cascade learning system in handling datasets with significant class imbalance. Most models may face a trade-off between accuracy and recall due to the imbalanced nature of HateSpeech. However, the accuracy-cost and recall-cost trade-offs, respectively depicted in the upper and lower subplots of Figure 3 and 4, demonstrate that our system effectively improves recall with minimal impact on accuracy as the cost budget increases. Although the recall rate of online cascade learning is marginally lower than the baselines under certain budgets, it can achieve a better balance between recall and precision, as evidenced by its consistently higher F1 scores in Appendix Figure 10 and 11. In particular, as demonstrated in Table 1 where Llama 2 70B Chat is the LLM and the cost budget $\mathcal{N}=4900$, our system even outperforms the LLM with a similar recall (ours: 82.03% vs. LLM: 82.19%) and a better accuracy (78.32% vs. 77.81%), underlining its effectiveness in handling imbalanced data streams.
@@ 5.3 Adaptability to Larger Cascade
To further validate the adaptability of our online cascade learning system, we have also explored scaling up the cascade system by further integrating a BERT-large model (4 models in total). The results of this integration are encouraging, detailed in Appendix Figure 12 and 13. The expanded cascade system, particularly when trained on Llama 2 70B Chat annotations, demonstrates equivalent or even superior results compared to the standalone LLM in most cases, with the exception of the HateSpeech dataset, a simpler task setting where a larger cascade might complicate the deferral policy learning and thus degrading overall performance. Therefore, it is important to align cascade size with task complexity to optimize performance and avoid overfitting. A more detailed analysis is in Appendix C.2
@@ Appendix A Detailed Proofs and Theoretical Analysis
This appendix section provides detailed proofs and explanations for the theoretical analysis presented in the main paper Section 3.1.
@@
$\gamma$ $=J(\pi)-\min\limits_{\pi\in\Pi}J(\pi)$ (8)
$=\sum^{T}_{t=1}\sum^{N}_{i=1}p^{s_{i}}_{\pi}C_{\pi}(s_{i})-\min\limits_{\pi\in\Pi}\sum^{T}_{t=1}\sum^{N}_{i=1}p^{s_{i}}_{\pi}C_{\pi}(s_{i})$ (9)
We simplify by defining the best fixed policy as $\pi^{*}=\operatorname*{arg\,min}_{\pi\in\Pi}J(\pi)$, leading to the following representation of regret:
@@
$=\sum^{T}_{t=1}\sum^{N}_{i=1}\prod^{i-1}_{j=1}p(\pi,s_{j})^{\prime}\cdot\Big{(}(1-p(\pi,s_{i})^{\prime})\cdot\mathcal{L}(a_{i}|x_{t})+p(\pi,s_{i})^{\prime}\cdot\mu c_{i+1}\Big{)}$
$\ \ -\sum^{T}_{t=1}\sum^{N}_{i=1}\prod^{i-1}_{j=1}p(\pi^{*},s_{j})^{\prime}\cdot\Big{(}(1-p(\pi^{*},s_{i})^{\prime})\cdot\mathcal{L}(a^{*}_{i}|x_{t})+p(\pi^{*},s_{i})^{\prime}\cdot\mu c_{i+1}\Big{)}.$ (11)
@@
To contextualize our proposed online cascade learning algorithm, we first analyze a simplified online ensemble learning algorithm. It consists of a sequence of models $m_{1},...,m_{N}$, each with a fixed operating probability $\sum_{i=1}^{N}w_{i}=1$ for prediction. Assuming a convex and differentiable cost function $c^{t}$ for all $t$, and that each model space $||M_{i}||$ is bounded, closed, and nonempty for all $i$, we define $||x||=\sqrt{x\cdot x}$ and $d(x,y)=||x-y||$ to establish:
@@
$\ \ \ \ \ \sum^{T}_{t=1}\sum^{N}_{i=1}\prod^{i-1}_{j=1}p(\pi,s_{j})^{\prime}\cdot\Big{(}(1-p(\pi,s_{i})^{\prime})\cdot\mathcal{L}(a_{i}|x_{t})+p(\pi,s_{i})^{\prime}\cdot\mu c_{i+1}\Big{)}$ (36)
$=\sum^{T}_{t=1}\sum^{N}_{i=1}\prod^{i-1}_{j=1}p(\pi,s_{j})^{\prime}\cdot\Big{(}p(\pi,s_{i})^{\prime}\big{(}\mu c_{i+1}-\mathcal{L}(a_{i}|x_{t}))+\mathcal{L}(a_{i}|x_{t})\Big{)}.$ (37)
Given that $p(\pi,s_{i})^{\prime}$ lies within the range $(0,1)$ for all $t$ and $i$, the coefficient preceding $\mathcal{L}(a_{i}|x_{t})$, namely $\prod^{i-1}_{j=1}p(\pi,s_{j})^{\prime}(1-p(\pi,s_{i})^{\prime})$, also falls within $(0,1)$. Assuming $\mathcal{L}$ is convex, and following the convergence arguments in Li & Orabona (2019), we establish that as $T\rightarrow\infty$, $\mathcal{L}(a_{i}|x_{t})$ converges to a minimal loss value $\epsilon_{i}$ for each model.
@@ B.1 A Simple Prefill Experiment
Figure 10: A more detailed cost-performance trade-off plot with accuracy, F1-score, recall, and precision curves, using GPT-3.5 Turbo as the LLM.
Figure 11: A more detailed cost-performance trade-off plot with accuracy, F1-score, recall, and precision curves, using Llama 2 70B Chat as the LLM.
Figure 12: Accuracy curve (and Recall curve for HateSpeech) with respect to costs, using GPT-3.5 Turbo as the LLM in a larger cascade that also comprises logistic regression, BERT-base, and BERT-large.
Figure 13: Accuracy curve (and Recall curve for HateSpeech) with respect to costs, using Llama 2 70B Chat as the LLM in a larger cascade that also comprises logistic regression, BERT-base, and BERT-large.
@@ B.3 Experimental Configurations
The detailed hyperparameter settings for online cascade learning are listed in Table 3 and 4. We tuned the hyperparameters using a grid search method on a separate validation set, which is a standard practice to avoid overfitting. In particular, we used the training set prepared for the knowledge distillation (as mentioned in Section 4.2) as our online cascade learning method’s validation set.
@@
Note that the “learning rate” in the table refers to the learning rates of the MLPs (in Section 3.2), not the models’. We used a consistent configuration for both online cascade learning and the distillation baselines by setting BERT-base’s batch size to 8, the learning rate to 0.00001, and the number of epochs to 5.
@@ C.1 Analysis of Training & Inference Cost Equilibrium
We conduct a thorough analysis of the training overhead for the models involved in our cascade to quantify the computational costs incurred in our approach. Below are the computation costs for training or inference over one sample (the computational costs of the confidence calibration MLP in Section 3.2, inference: $897\ Flops$, training: $1794\ Flops$, are negligible):

### v3

@@ 4 Experimental Setup
We construct a cascade system using three models: (i) a logistic regression model, (ii) a pretrained BERT-base model with 110M parameters (Kenton & Toutanova, 2019), and (iii) GPT-3.5 Turbo [^1] https://platform.openai.com/docs/models/gpt-3-5. This diverse set of models allows us to evaluate the effectiveness and robustness of our approach across various complexities and types of queries.
@@
Benchmarks
We evaluate our approach on four benchmarks that reflect the demands of real-world streaming applications in a wide range of commercial services, from customer feedback analysis to content moderation:
  • IMDB. A binary sentiment classification benchmark with 50,000 movie reviews (Maas et al., 2011). The dataset has an even distribution of positive and negative samples. We use the official training split for our experiments, which contains 25,000 samples.
  • HateSpeech. A binary classification dataset consisting of posts from an online forum, annotated with hate and noHate labels (de Gibert et al., 2018). After filtering, the dataset contains 10,703 samples with a pronounced class imbalance (1:7.95 ratio) between hatespeech and non-hatespeech examples. This imbalance mirrors a realistic challenge in streaming data environments, particularly in detecting harmful content. Our evaluation on this benchmark focuses on both accuracy and recall.
  • ISEAR. A multi-class emotion detection benchmark encompassing 7,666 samples across seven categories (Joy, Fear, Anger, Sadness, Disgust, Shame, Guilt) (Shao et al., 2015). Each category is well-represented, providing a balanced label distribution across the dataset.
  • FEVER. A fact-checking dataset with 6,512 claims manually verified against Wikipedia, labeled as Supported or Refuted (Thorne et al., 2018). It tests our framework’s ability to perform complex reasoning and information verification, a crucial aspect for real-time truth assessment in streaming data applications.
@@
Baselines
We compare online cascade learning against several baselines to establish its effectiveness:
@@
  • Online Ensemble Learning: We employ all available models in an ensemble with learned predetermined probabilities. The smaller models are also continuously updated based on LLM annotations. This serves as an ablation of our method by excluding the deferral policy learning component.
@@ 5.1 Overall Performance and Cost Trade-offs
The results on the HateSpeech dataset further reveal the strengths of our online cascade learning system in handling datasets with significant class imbalance. Most models may face a trade-off between accuracy and recall due to the imbalanced nature of HateSpeech. However, the accuracy-cost and recall-cost trade-offs, respectively depicted in the upper and lower subplots of Figure 3 and 4, demonstrate that our system effectively improves recall with minimal impact on accuracy as the cost budget increases. Although the recall rate of online cascade learning is marginally lower than the baselines under certain budgets, it can achieve a better balance between recall and precision, as evidenced by its consistently higher F1 scores in Appendix Figure 10. In particular, as demonstrated in Table 1 where Llama 2 70B Chat is the LLM and the cost budget $\mathcal{N}=4900$, our system even outperforms the LLM with a similar recall (ours: 82.03% vs. LLM: 82.19%) and a better accuracy (78.32% vs. 77.81%), underlining its effectiveness in handling imbalanced data streams.
@@ 5.3 Adaptability to Larger Cascade
To further validate the adaptability of our online cascade learning system, we have also explored scaling up the cascade system by further integrating a BERT-large model (4 models in total). The results of this integration are encouraging, detailed in Appendix Figure 11. The expanded cascade system, particularly when trained on Llama 2 70B Chat annotations, demonstrates equivalent or even superior results compared to the standalone LLM in most cases, with the exception of the HateSpeech dataset, a simpler task setting where a larger cascade might complicate the deferral policy learning and thus degrading overall performance. Therefore, it is important to align cascade size with task complexity to optimize performance and avoid overfitting. A more detailed analysis is in Appendix C.2
@@ Appendix A Detailed Proofs and Theoretical Analysis
This appendix section provides detailed proofs and explanations for the theoretical analysis presented in the main paper Section 3.
@@
$\gamma$ $=J(\pi,T)-\min\limits_{\pi\in\Pi}J(\pi,T)$ (6)
$=\sum^{T}_{t=1}\sum^{N}_{i=1}p^{s_{t,i}}_{\pi}C_{\pi}(s_{t,i})-\min\limits_{\pi\in\Pi}\sum^{T}_{t=1}\sum^{N}_{i=1}p^{s_{t,i}}_{\pi}C_{\pi}(s_{t,i})$ (7)
As the value of $t$ is clear from the context, we abbreviate states $s_{t,i}=\langle x_{t},i\rangle$ by $s_{i}$. We denote the best fixed policy as $\pi^{*}=\operatorname*{arg\,min}_{\pi\in\Pi}J(\pi,T)$ and $\pi(s_{t,i},\texttt{defer})$ as $p(\pi,s_{i})^{\prime}$. Furthermore, since the prediction loss is computed as the aggregation over a class probability distribution, we simplify $\sum_{y\in Y}\pi(s_{t,i},y)\cdot\mathcal{L}(y|y_{t})$ as $\mathcal{L}(a_{i}|y_{t})$ by using $a_{i}$ to represent the output probability vector, which leads to the following representation of regret:
@@
$=\sum^{T}_{t=1}\sum^{N}_{i=1}\prod^{i-1}_{j=1}p(\pi,s_{j})^{\prime}\cdot\Big{(}(1-p(\pi,s_{i})^{\prime})\cdot\mathcal{L}(a_{i}|y_{t})+p(\pi,s_{i})^{\prime}\cdot\mu c_{i+1}\Big{)}$
$\ \ -\sum^{T}_{t=1}\sum^{N}_{i=1}\prod^{i-1}_{j=1}p(\pi^{*},s_{j})^{\prime}\cdot\Big{(}(1-p(\pi^{*},s_{i})^{\prime})\cdot\mathcal{L}(a^{*}_{i}|y_{t})+p(\pi^{*},s_{i})^{\prime}\cdot\mu c_{i+1}\Big{)}.$ (9)
@@
To contextualize the no-regret analysis of our proposed online cascade learning algorithm, we first analyze a simplified online ensemble learning algorithm under the same stream processing setting that comprises the linear combination of a series of classification models $m_{1},...,m_{N}$, each with a static operating probability $\sum_{i=1}^{N}w_{i}=1$. Let us denote the model parameters of $m_{i}$ at time $t$ by $m_{i}^{t}$. Assuming a convex, differentiable cost function $c^{t}$ for all $t$ that can evaluate $m_{i}^{t}$, and bounded, closed, nonempty model spaces $||M_{i}||$ for all $m_{i}$, we define $||x||=\sqrt{x\cdot x}$ and $d(x,y)=||x-y||$ to establish:
@@
$\ \ \ \ \ \sum^{T}_{t=1}\sum^{N}_{i=1}\prod^{i-1}_{j=1}p(\pi,s_{j})^{\prime}\cdot\Big{(}(1-p(\pi,s_{i})^{\prime})\cdot\mathcal{L}(a_{i}|y_{t})+p(\pi,s_{i})^{\prime}\cdot\mu c_{i+1}\Big{)}$ (34)
$=\sum^{T}_{t=1}\sum^{N}_{i=1}\prod^{i-1}_{j=1}p(\pi,s_{j})^{\prime}\cdot\Big{(}p(\pi,s_{i})^{\prime}\big{(}\mu c_{i+1}-\mathcal{L}(a_{i}|y_{t}))+\mathcal{L}(a_{i}|y_{t})\Big{)}.$ (35)
Given that $p(\pi,s_{i})^{\prime}$ lies within the range $(0,1)$ for all $t$ and $i$, the coefficient preceding $\mathcal{L}(a_{i}|y_{t})$, namely $\prod^{i-1}_{j=1}p(\pi,s_{j})^{\prime}(1-p(\pi,s_{i})^{\prime})$, also falls within $(0,1)$. Assuming $\mathcal{L}$ is convex, and following the convergence arguments in Li & Orabona (2019), we establish that as $T\rightarrow\infty$, $\mathcal{L}(a_{i}|y_{t})$ converges to a minimal loss value $\epsilon_{i}$ for each model.
@@ B.2 LLM Prompts
Figure 10: A more detailed cost-performance trade-off plot with accuracy, F1-score, recall, and precision curves, respectively using GPT-3.5 Turbo and Llama 2 70B Chat as the LLM.
Figure 11: Accuracy curve (and Recall curve for HateSpeech) with respect to costs, respectively using GPT-3.5 Turbo and Llama 2 70B Chat as the LLM in a larger cascade that also comprises logistic regression, BERT-base, and BERT-large.
@@ B.3 Experimental Configurations
The detailed hyperparameter settings for online cascade learning are listed in Table 3 and 4. We tuned the hyperparameters using a grid search method on a separate validation set, which is a standard practice to avoid overfitting. In particular, we used the training set prepared for the knowledge distillation (as mentioned in Section 4) as our online cascade learning method’s validation set.
@@
Note that the “learning rate” in the table refers to the learning rates of the MLPs (in Section 3: Confidence Calibration), not the models’. We used a consistent configuration for both online cascade learning and the distillation baselines by setting BERT-base’s batch size to 8, the learning rate to 0.00001, and the number of epochs to 5.
@@ C.1 Analysis of Training & Inference Cost Equilibrium
We conduct a thorough analysis of the training overhead for the models involved in our cascade to quantify the computational costs incurred in our approach. Below are the computation costs for training or inference over one sample (the computational costs of the confidence calibration MLP in Section 3, inference: $897\ Flops$, training: $1794\ Flops$, are negligible):